OpenStudy (anonymous):

Find the exact value of the expression. tan(sin^-1(4/5))

OpenStudy (anonymous):

\[\tan (\sin^{-1} (4/5))\]

OpenStudy (anonymous):

Is the correct answer 4/3?

OpenStudy (mertsj):



OpenStudy (mertsj):

You can see that it is.

OpenStudy (mertsj):

tan A = 4/3 and sin A = 4/5

OpenStudy (mertsj):

Vengeance is incorrect

OpenStudy (anonymous):

Ok, thank you

OpenStudy (anonymous):

Ah yes sorry I had the calculator in the RAD mode.